OpenStudy (anonymous):

Solve 0.25(7 + 3m) = -0.125 m.

OpenStudy (anonymous):

0.25(7 + 3m) = -0.125 m distribute 0.25(7 + 3m) first.

OpenStudy (anonymous):

1.75+0.75m

OpenStudy (anonymous):

0.25(7 + 3m) = -0.125 m 1.75+0.75m = -0.125m 0.75m = -1.875 m = -2.5